# The Geometry behind the Two-Dimensional Hard-Disk ‘Glass Transition’

### v1

@@ VIII.1 Introducing the Floret Pentagonal Tiling
The structure in Fig. 6b is known from the field of pentagonal tilings, i.e., it is a monohedral, convex, pentagonal tiling of type 6 [72, 73]. It is also often referred to as a floret pentagonal tiling (FPT). I will refer to the particle stacking into ordered wheels as an FPT structure in recognition of this connection. For my intents and purposes, however, it is important that the tiling follows from a Voronoi tessellation argument that respects the size of the particles and has congruent tiles.
Note that the Voronoi vertex in the center of each wheel is 4-fold degenerate. That is, if one were to consider a wheel in isolation and perturb the particle positions from their contact, one would be able to gain 3 additional vertices by some of the particles becoming neighbors. This could be done, e.g., by reducing the overall area fraction, which would give each particle freedom to move. In practice, there are thus 6 Voronoi-cell edges per particle on average in the FTP structure. This is likely why it is so difficult to pick up on FTP Voronoi-cell signals in the thermalized systems.
VIII.2 A Geometric Ground State
It is clear that the structure in Fig. 6b is a crystalline solid and mechanically rigid. In particular, it cannot be melted into a fluid at a constant $\phi$. In fact, one can remove particles from each wheel, without impairing the rigidity, as each particle in the system has 5 contact points. At most two particles per wheel can be removed, for a structure that is still ordered at a longer range and maintains rigidity. This does come at the cost of having two types of Voronoi cell, which I consider undesirable, as will become clear in Section VIII.5. The FPT has a single realization and this means its configurational entropy is zero. Removing or adding particles moves $\phi$ (slightly) away from $\sqrt{3}\pi/7$ in either direction. However, it simultaneously strongly increases the configurational entropy. For a monodisperse hard-disk system, I will refer to the FPT as a ‘geometric ground state’ at $\phi=\phi_{g}$i.e., it is the lowest entropy configuration in the ($T\downarrow 0$) limit.
Now we can turn to bidisperse (possibly also polydisperse) systems. In the thermodynamic limit ($N\uparrow\infty$), one can easily discount edge effects and place all particles of the same size in their own FPT. That is, one would have several contacting areas of specific size particles in FTP with mismatching lattice size. Thus, provided features of the structure in Fig. 6b are responsible for the appearance of $\phi_{a}$, the associated $\phi$ should be (relatively) independent of $R$. This agrees with my findings for $R\lesssim 0.8$ in Section VI, and would also explain the appearance of some structural features at $\phi\approx 0.777$ for $R\gtrsim 0.8$. Any variations that were observed on these are likely due to finite-size effects or other physical mechanisms interfering, such as proximity to the ‘critical point’ in $n_{\bar{q}}=5.5$. Turning to Appendix A, it then also becomes clear why point “g” in Fig. 8, which was obtained for another stoichiometric ratio (1:2) [96], is close to $\phi_{g}$. In principle, if there is only a geometric requirement, then it is likely that signatures of $\phi_{g}$ can be recovered in polydisperse systems, provided there is sufficient statistics.
VIII.3 Glass, Frustrated Disorder, or Something Else
Systems with $R\approx 1$ tend to crystallize into a hexagonal configuration, rather than an FTP structure. This is well-known from the literature [96] and makes sense. That is, intuitively the disks can make their density homogeneous by locally arranging into a hexagonal crystal rather than an FPT. Thermodynamically, the state in Fig. 6b is also unfavored for $T>0$, because there is only a single realization, for which all Voronoi cells are identical and pentagonal. When assuming a hexagonal crystal configuration at $\phi=\phi_{g}$, the particles can vibrate about their ideal hexagonal-lattice position, raising both the configurational and vibrational entropy of the system.
The shape of the red curve in Fig. 4 suggests that the role of $R$ is akin to that of temperature in a regular phase diagram. This led me to adopt the wording ‘critical point,’ when referencing the left-most point of the red curve in the state diagram, see Section VI. Given that $\phi_{a}$ is extracted from the limiting behavior of the diffusition coefficient $D(\phi)\downarrow 0$, it is sensible to assume that this is not a critical point in the equilibrium sense of the word. Note, however, that while the long-time linear part to the MSD is quickly dropping off, there is displacement at sufficiently long times. Thus, to state definitively that the system has fallen out of equilibrium is not accurate.
This is also evidenced by the SISF data in the experiment of Lozano et al. [1], which only shows a divergent relaxation time near rcp. I therefore conclude that there is no glassy dynamics at their measured $\phi\approx 0.776$ and that the (nearly) constant feature at $\phi\approx 0.777$ in my simulations is present in equilibrium. I used the time-dependent data to gain additional insight by analyzing the way neighborhoods change, see the ESI [59].
Figure 7: The relative neighborhood exchange rate indicates strong fluctuations close to transitions. The ratio of the number of changes between a pentagonal and a hexagonal neighborhood to the total number of particle displacements in the system, $r_{5\leftrightarrow 6}$, as a function of the area fraction $\phi$. The red and blue points show data gathered for $R=1$ and $R^{-1}=1.4$, respectively, as labelled. The solid, light-red and light-blue curves serve to guide the eye. The peak positions are indicated by the magenta and purple vertical dashed lines, respectively. Errors are smaller than the symbol size and therefore not shown here.
In brief, for each time step, I compute the radical Voronoi diagram and keep track of the $q$ value belonging to each particle. I find that for sufficiently high $\phi$, the neighborhoods dwell about integer values of $n_{q}$ and the system does not favor half-integer $n_{q}$. For example, the trajectory of a certain particle may lead it to explore pentagonal ($n_{q}\approx 5$) and hexagonal ($n_{q}\approx 6$) neighborhoods, but spend little time in $5.5$-gonal neighborhoods  [^7] This observation does not conflict with the fact that in Fig. 2a-c, I observe dips in the $P_{\phi}(q)$ about integer values of $n_{q}$ and a smooth increasing trend near half-integer values of $n_{q}$. There is a bias to the representation of the system using $P_{\phi}(q)$, as explained in Section IV.. Next, I extract transition rates for the exchange between neighborhoods and normalized this by the total number of particle displacements in the system. This resulted in, for example, relative exchange rates between pentagonal and hexagonal neighborhoods, $r_{5\leftrightarrow 6}$, some which are shown in Fig. 7.
Surprisingly, the relative number of exchanges constitutes a substantial fraction of particle displacements even at high $\phi$ close to $\phi_{a}$ or equivalently $\phi_{g}$. This means that while there is relatively limited cage hopping, as evidenced by the low value of $D(\phi)$, the cages themselves are quite dynamic. Additionally, I find no significant difference in the $r_{i\leftrightarrow i+1}$ between small and large particles for any integer $i$, see the ESI [59]. This implies that the cage dynamics is shared equally among the large and small particles, though it does not exclude a difference in cage-hopping rate, i.e., in diffusivity.
Both data sets presented in Fig. 7 show a peak. When there is crystallization, as is the case for $R=1$, it is sharp and very pronounced. Fitting the peak values results in $\phi\approx 0.703$ and $\phi\approx 0.757$ for $R=1$ and $R^{-1}=1.4$, respectively. The former agrees very well with where the fluid state is expected to become unstable. The latter is close to, but not quite at $\phi_{g}$. For $R^{-1}=1.4$, the peak is also broader and less pronounced. Within the error, I detected no discernable drift toward one type of neighborhood as a function of time. This suggests that there is no slow relaxation toward a phase-separated state, as would be the case for glassy dynamics.
That the rate of cage rearrangement is relatively high, fits with the conceptual picture of an active probe particle (experiments in Ref. [1, 61]) picking up on the fluctuations of the local environment, as put forward in recent modeling work from my group [60, 61]. The fact that $r_{5\leftrightarrow 6}$ is asymmetric about $\phi\approx\phi_{g}$ would also fit with the asymmetric probe response observed in Ref. [1]. However, there is a slight mismatch between the peak value of $\phi$ and $\phi_{g}$. This should be addressed in future work, wherein active probes are included in the sample, but this falls outside of the scope of the present research.
Weighting the evidence, I believe that the frustration experienced by the system at $\phi=\phi_{g}$ can be attributed to the special properties of the FTP lattice, rather than kinetic arrest. That being said, the geometric ground state may be a deep (local) entropic minimum for $T=0$, but at finite temperature, it is not obvious why the FPT should be important. In fact, it is extremely likely that FTP cannot form for $T>0$. I will address how to reconcile these two seemingly conflicting points in Section VIII.5.
VIII.4 Other Geometric Ground States
Before getting to that, let us consider if there is additional evidence that could support my viewpoint by examining other geometric-ground-state candidates. An equilateral-triangle tessellation of Voronoi cells forms when the particles are placed in contact in a regular honeycomb lattice. Each particle then has 3 neighbors, which makes the structure mechanically unstable, unless the particles are fully frictional. The close-packing honeycomb lattice gives an area fraction $\phi_{t}=\pi/(3\sqrt{3})\approx 0.605$. This is the lowest density lattice, for which the particles can be made to touch in such a way that equal-shape triangular Voronoi cells are obtained.
Note that $\phi_{t}$ is close to where I observe an inflection point in the derivative of the diffusion coefficient, see Section V. In fact, the guide to the eye in Fig. 5d is a fitted constant with value $\phi\approx 0.607$ to all the data points, which matches the $\phi_{t}$ value very well. In addition, the predicted honeycomb-lattice $\phi$ is close to where Li et al. [2] locate their “onset glass transition” / the appearance of caging in the fluid. Adopting the geometric-ground-state viewpoint, it makes sense that for the $\phi=\phi_{t}$, the response of the fluid to external perturbations is strongest.
Turning to regular, square Voronoi cells, I find that they form by a simple-square stacking of disks with $\phi=\pi/4$. This structure is isostatic, but not stable to shear deformation for frictionless disks. The $\phi$ value lies close to, but measurably above, where I report $\phi_{a}$ in the system. There are several potential signals of a square-only ground state. Firstly, the area fraction associated with the ‘critical point’ in $n_{\bar{q}}=5.5$ curve: $\phi\approx 0.785$ for $R^{-1}\approx 1.21$. This value matches closely $\phi=\pi/4$, although I acknowledge there is substantial uncertainty to determining this point. Potentially, the $P^{\ast}_{5}$-minimum data in Fig. 5d can also be ascribed to the presence of a square-tiling ground state. However, the mean value of this data appears to lie at a slightly higher value of $\phi\approx 0.793$. This is both within two standard deviations ($\sigma_{\phi}\approx 0.005$) of the $\phi=\pi/4$ and the $\max_{\phi}\!\left(n_{\bar{q}}\right)$ line. It is therefore difficult to draw any substantial conclusions about the effect of a potential square-tiling geometric ground state. Presumably this is also because $\phi_{g}<\pi/4$ and associated with a much more stable FTP tiling.
VIII.5 Connection to Foam-like Tissue Models
I had set out to make the connection to the literature on foam-like tissue modelling [62, 63, 64, 65, 66, 67, 68, 69]. Taking on this viewpoint, one can see the presence of the FPT ground state at $\phi=\phi_{g}$ as the particle equivalent to imposing a shape index belonging to a regular pentagon for epithelia/foam models. If this is a correct viewpoint, a geometric or topological reasoning may be given for the appearance of the structural and dynamic features at $\phi=\phi_{g}$.
The first thing to note is that the connection is not simple. The value of $q$ associated with the FTP lattice is $q\approx 0.777$, which is close to $q_{\mathrm{r}}(4)=\pi/4\approx 0.785$. In a foam-like model with a Hamiltonian of the form in Eq. (1), the equivalent shape index would place the system firmly in the fluid-like state. To observe jamming in such a system requires $q\approx 0.865$ (value for a regular pentagon) [62]. This suggests that a direct mapping to foam-like models with Eq. (1) is not possible.
Nonetheless, there are several aspects that suggest there is a path forward. Examining the single-particle neighborhood in my granocentric analysis, I found that the isoperimetric quotient $q$ is identical to the associated (local) volume fraction $\phi_{g}$. This seems to be a robust result for a range of neighborhood configurations involving particles that are in contact, see the ESI [59]. Contact-preserving distortions may therefore leave $\bar{q}$ unaffected. Note, however, that this constancy does not hold when the particles are out of contact. That is, the observed mean value of $q$ is not the same as the FTP one. In the range where $\phi_{a}$ is nearly constant, I measure $\bar{q}\approx 0.88$. This need not immediately discount the intuition about the geometric ground state. For the foam-like tissue models, the mean shape index is also not necessarily equal to the one that is imposed.
Nonetheless, it is well known that certain topological properties are invariant under diffeomorphism, as are other invariants [98]. I have already stated that the FTP structure cannot be broken up to form a fluid (there are 5 neighbors per particle), nor can it be formed from a fluid phase in a defect-free manner. However, consider a FTP structure which has two adjacent vacancies in one of its wheels, but is otherwise perfect. Then it is clear that the structure may be systematically ‘stripped’ from this central point out into a crystalline structure, which may subsequently be made (to appear) disordered. Thus it seems likely that a diffeomorphism, barring an intensive subset of the FTP lattice, can be constructed.
At present, a full mathematical framework by which the connection can be proven, appears out of reach. Other candidates for making the connection involve the study of cage transformations and T1-transitions [62, 99, 100, 101, 102, 103, 104]. It may also be possible to repurpose the quasi-species approach of Ref. [105], as I suspect the behavior that I find is in equilibrium. However, for now, such an effort must be left to future work.
@@ IX Discussion
Thus far, I have avoided speaking of the ‘ideal glass transition’. That is, the presence of a disordered system of particles with zero configurational entropy. This has been a topic of significant interest [106, 96, 107], with a very recent study focusing on 2D [108]. Given the requirements for a glass transition, the constant $\phi_{a}$ range cannot be identified as such, as I observed earlier in Section VIII.3. Nonetheless, there is compelling evidence for a structural transition that underlies my observations. This transition I would identify as a continuous one, given the smoothness of the $R\lesssim 0.83$ curves in Appendix D.
I would further like to note that the concept of a geometric reference is by no means a new one. However, it is typically applied locally, for example, when examing the dynamics in glasses [22, 9, 23, 2, 49, 42, 109]. What has not been previously reported, to the best of my understanding, is that seemingly disordered (2D hard-disk) particle fluids can become entangled with geometric, regularly ordered ground states. Given the recent successes of machine learning techniques in determining the dynamic propensity of particles [36, 37, 40], it seems likely that the application of such a machine-learning approach would be able to pick up on $\phi_{a}$.
Thus far, I have spoken about geometry, which makes sense for hard-particle systems. In systems where the particles interact via soft potentials, it is not obvious whether the argument can be made by solely by considering the particle arrangement. However, I would expect that elements of the geometric ground state are preserved in such soft systems. I base this intuition on fact that foam-like tissue models [62, 63, 64, 65, 66, 67, 68, 69] have Harmonic potentials governing their dynamics, yet show jamming due to geometric frustration.
I alluded to the potential for random-loose packing (rlp) to be associated with the appearance of glassy dynamics in my comparison to the literature in Appendix A. Reflecting on this, I would think that this relation is tenuous at best. The FTP has 5 neighbor contacts, which is well above the isostatic requirement of $3$ (frictional) or $4$ (frictionless) contacts. I suspect that the origin of the observation in Ref. [110] may be attributed to the intervention of the transition that I have picked up in the data. That is, the presence of the FTP geometric ground state might have biased the simulation toward placing rlp close to $\phi_{g}$. The rlp connection in the numerical work of Refs. [111, 112] could be similar in nature. However, the connection is also found in 3D systems.
It is possible that the notion of a geometric ground state can be transferred to 3D systems. In Appendix F, I analyze several candidate structures. In foam-like tissue modeling, there is evidence that a geometric frustration similar to the one in 2D can occur [113]. In addition, when the particles are arranged using the underlying structure of a (non-gyrated) tetrahedral-octahedral honeycomb, I obtain a volume fraction of $\eta=8\pi/(27\sqrt{3})\approx 0.537422$. In this arrangement, the particles are the inscribed spheres to the octahedra comprising the honeycomb. This packing fraction is relatively close to the reported rlp value for frictional spheres in 3D, $\eta\approx 0.536$ [44], though it is often placed at a slightly higher value of $\eta\approx 0.555$ in experimental studies [114]. However, it should also be clear that 3D systems are quite different from their 2D counterparts. The exploration of 3D is therefore left to follow-up work.
Lastly, I should address potential for experimental validation. Local rheological characterization using active particles [1] and perturbing the system via local laser illumination [2], both have the potential to reveal structural and dynamical changes. It would be useful to see to what extent either approach reveals the presence of near-constant $\phi$ features in the $R-\phi$ state diagram, as predicted by the simulations performed here. It would additionally be useful to identify the presence of the underlying FTP structure directly. However, at present it is less clear how to approach this without gathering considerable statistics, as was done here. Future work will focus on developing the theory behind the connection and this effort may reveal a more direct quantifier.

### v2

@@ VIII.1 The Floret Pentagonal Tiling
The structure in Fig. 8b is known from the field of pentagonal tilings, i.e., it is a monohedral, convex, pentagonal tiling of type 6 [104, 105]. It is also often referred to as a floret pentagonal tiling (FPT). I will refer to the particle stacking into ordered wheels as an FPT structure in recognition of this connection.
VIII.2 Geometric Ground States
The FPT structure can also be seen as the result of T2 removals of one in seven of the particles in a hexagonal close packed arrangement, see Fig. 2b. In particular, one where the removals lead to a regular sublattice of voids. Interestingly, starting from a crystalline solid (at infinite pressure), this is the first (lower) area fraction achieved by T2 removals, for which both the voids are arranged regularly, and all neighborhoods are identical. Equivalently, all Voronoi tiles are congruent and their edges are by construction tangent to the inscribed sphere. This is important, since — as mentioned previously — for a monodisperse system where the disks are inscribed to a congruent (tangential) tiling, the equivalence $\phi=q$ holds. That is, the area fraction obtained from the MSD and shell-percolation criterion is an indicator of the relevance of the Voronoi cell belonging to this tiling.
A second such configuration exists, where two out of three particles are removed, with the remaining ones forming a regular honeycomb lattice. This corresponds to a Voronoi tessellation where each cell is congruent and an equilateral triangle. The associated area fraction $\phi=\pi/(3\sqrt{3})\approx 0.6046$, very close to our fitted constant value of $\phi_{i}\approx 0.607$ for the inflection point to $\partial D/\partial\phi$. Note that this is removed from where the monodisperse system has liquid-hexatic coexistence ($\phi\approx 0.699$ for the liquid and $\phi\approx 0.716$ for the hexatic phase coexistence area fractions), as well as the continuous hexatic-solid transition that is located at $\phi\approx 0.724$ [126, 127, 121, 122, 128]. Therefore, it is reasonable to say that the transition related to the inflection point is not related to the hexatic phase transition.
For the FPT-based transition there may be a relation to the hexatic phase transition. Above, I related the FPT structure to T2 removals, which can form pairs of pentagons and heptagons in a hexatic lattice. This kind of pair formation is also reported in defect unbinding for the hexatic phase transition [121, 128, 94]. In addition, the diffusivity data $\phi_{a}$ runs from $\approx 0.7$ to the value $\phi\approx 0.777$ in a(n apparently) smooth fashion, see Fig. 6, given the quality of the data. Thus, the FPT-based transition, which I claim to underlie the appearance of arrested dynamics, may be a limiting case of the topological defect unbinding. This connection will need to be analysed further in future work.
Both the FPT and honeycomb correspondence suggest that there is merit to identifying regular geometric ground states that result from T2 removal of particles from the close-packed structure and possess congruent Voronoi tiles. It is possible to form other structures with congruent Voronoi tessellations, but these do not arise naturally from T2 transitions of the close-packed crystalline structure and are therefore not considered here. I revisit this possibility again in 3D in the Discussion section, but there too the most promising connection follows by considering T2 particle removals starting with the closest packed structure.
Note that at zero temperature (infinite pressure) the structure in Fig. 8b is a crystalline solid and mechanically rigid. This begs the question of how these structures influence clearly disordered states of bidisperse particles. I hypothesize that transforming a geometric ground state can be accomplished by bringing together two bulk FPT states and introducing a few defects into the structure. These defects allow for freedom of motion and a systematic reshaping of the structure to something that looks otherwise like an amorphous system. Alternatively, the Mermin-Wagner theorem — forbidding long-ranged order in 2D — could provide the necessary fluctuations to destroy the FPT order and leave an amorphous state at finite pressure (and temperature). The influence of the state could then be explained by the topological nature of the transitions required to rearrange particles, neighbor exchanges involve T1 transitions not T2 transitions. It is well known that topological properties are invariant under diffeomorphisms, as are other invariants [129]. The mechanism for the particle system would be similar to the way the isoperimetric target value of the Hamiltonian (1) controls the dynamics of a model foam or tissue, as $p_{0}$ is not influenced by neighbor exchanges.
For the honeycomb lattice, there is no intrinsic rigidity (save for fully frictional particles) and smoothly distorting this configuration into that of an (arrested) amorphous fluid phase is easier to understand. That is, I expect that the value of $\phi_{i}$ for a bidisperse system can be explained by similarly bringing together two perfect honeycomb lattices with different particle sizes and allowing these to rearrange via thermal motion into a disordered fluid-like state. Unlike for the FPT state where $P_{5}^{\ast}$ showed a clear cusp, I found no hints of features in the local structure at $\phi\approx 0.607$ that could justify taking the limit $R\uparrow 1$ for this near-constant value of $\phi_{i}$. This makes the geometric undestanding more tenuous in for this value of $\phi$. I will address these issues next by performing additional computations.
VIII.3 Validation
Formalizing mathematically the above understanding is not trivial. For the foam and active tissue models, the connection to (soft glassy) rheology is made through the trap model [130, 86, 131]. But given the hard nature of the interaction in the bidisperse disk system investigated here, this may not be readily transferable. Mechanical topological features were only recently discussed in the context of hyperform materials [132]. However, it is possible to give an additional piece of information to support my viewpoint of geometric ground states.
If such a state is indeed responsible for the dynamic arrest, then the area fraction $\phi\approx 0.777$ should also appear from other compositions of the system. Turning to Appendix A, this may explain why point “h” in Fig. 10, which was obtained for another stoichiometric ratio (1:2) [133], falls within the range reported by Refs. [2, 3]. This connection can be made stronger by performing an additional numerical study on a sufficiently different system in terms of radii and composition.
Figure 9: Validating the geometric ground state Logarithmic plot of the reduced diffusion coefficient $D$ as a function of $\phi$ for a trinary with particle sizes $\sigma$, $0.8\sigma$, and $0.6\sigma$ in 1:1:1 stoichiometric ratio. The error is typically much smaller than the symbol size. The red curve shows a power-law fit to the data. The vertical dashed vertical magenta line indicates the inflection point to the change in diffusion. The inset shows the numerical derivative $\partial D/\partial\phi$ as a function of $\phi$. The solid orange curve shows a septic polynomial fit to this data and the magenta line is the same as in the main panel.
Figure 9 shows the diffusion coefficient obtained for a system of 2,400 hard disks with sizes $\sigma$, $0.8\sigma$, and $0.6\sigma$ in a 1:1:1 stoichiometric ratio. Data was processed in a manner analogous to that for our bidisperse system, although I equilibrated the system for longer, namely by $10^{5}$ diffusion times. I found the extrapolated dynamical-arrest point to be located at $\phi_{a}=0.775\pm 0.004$ and the inflection point at $\phi_{i}=0.59\pm 0.04$. Note that it remained difficult to localize the inflection point accurately. However, it is clear that the area fraction belonging to the floret pentagonal tiling was recovered, which I further confirmed by analyzing features of the shell-percolation transition, see the ESI [64].
@@ IX Discussion
Given the above results, there are several points worth critiquing. It is well known that systems with $R\approx 1$ tend to locally crystallize [133], rather than locally form an FPT structure  [^12] Technically the Mermin-Wagner theorem forbids this structure in 2D at finite temperature and pressure.. That is, the disks can make their density more homogeneous by locally arranging into a hexagonal crystalline arrangement rather than an FPT. Thermodynamically, the state in Fig. 8b is also unfavored for $T>0$, because there is only a single realization, for which all Voronoi cells are identical and pentagonal. When assuming a hexagonal crystal configuration at $\phi\approx 0.777$, the particles can vibrate about their ideal hexagonal-lattice position, raising both the configurational and vibrational entropy of the system.
Thus, it is difficult to imagine that a geometric ground state, as introduced here, would be able to influence (crystalline) configurations in a thermal or Brownian suspension. However, the numerical evidence presented here, combined with 2D results from other sources [133, 135, 57, 136, 137, 138], see Appendix A, in addition to the experimental work by Lozano et al. [1] and Li et al. [4] for Brownian systems, and that of Naseer et al. [2], Vishali et al. [3], and Puckett and Daniels [82] for granular disks, strongly suggests the following. The departure away from monodispersity (and potentially from frictionless interactions) allows for the geometric ground states that I report to compete with the driving forces that induce crystallization. This would explain why the indications for the presence of such states is so weak for nearly monodisperse systems, although there is circumstantial evidence in the simulation and theory literature [136, 137].
It would be interesting to see to what extent experimental evidence along the lines of the work carried out by Vishali et al. [3] or frictional simulations like those of Refs. [135, 81] are able to validate this viewpoint, as well as help identify topological features near $\phi\approx 0.777$. However, a force-network analysis and shear-based rheology constitute separate studies and are therefore left for future work. Narinder et al. [66] reported a correlation between the structure of a passive rod fluid and the rotational dynamics of an active particle that interacted with this fluid. A contact-frequency argument was made to explain this connection. In Appendix F, an analysis of the time-dependence of the change in neighborhoods is provided, also see the ESI [64], as a setup to such an analysis applied to the disk system studied here. This reveals that there are indeed rapid dynamics in the form of neighbor exchanges, which peaks (weakly) at a value close to $\phi_{g}$. This is a piece of evidence in support of the conceptual framework introduced in my group to explain the rotational diffusion enhancement for a self-propelled probe particle [65, 66]. Referencing Appendix F, the peak appears more pronounced for a monodisperse sample. Examining the properties of (quasi-)2D arrangements further using self-propelled probes or laser-type perturbations [4] might reveal more information, however, this too constitutes a separate project.
The concept of a geometric reference state is by no means a new one. Yet, it is typically applied locally, for example, when examining the dynamics in glasses [26, 11, 27, 4, 53, 45, 139]. What has not been previously reported, to the best of my understanding, is that seemingly disordered (2D hard-disk) particle fluids can become ‘topologically entangled’ with ordered, geometric ground states. Nonetheless, this is not a peculiar way of thinking about 2D systems, when referencing the literature on foam and tissue models [85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95]. It seems plausible that elements of 2D models based on geometric harmonic potentials be transferred in an effective manner to particles that interact via short-ranged (highly repulsive) potentials. This then suggests a possible common origin behind observations of a glass and jamming transition in 2D hard disks and model tissues, where the former is controlled by the FPT and the latter by an overlapping arrangement of regular pentagons. A potential route toward tackling this problem is to find a transformation that smoothly takes one proposed geometric ground state into another. I see projections of higher-dimensional regular lattices on a 2D subspace via a cut-and-project type operation [140, 141, 142] as a viable route of exploration in this direction. Speculating, the angle of the plane would have to be varied between rational and irrational to achieve the desired transformation.
Lastly, the notion of a geometric ground state may be transferred to 3D systems. In Appendix G, I analyze several candidate structures. In foam-like tissue modeling, there is evidence that a geometric frustration like the one in 2D can occur [143]. In 3D, the random loose packing (RLP) of spheres is reported for a volume fraction $\eta\approx 0.555$ in various studies [144, 135, 75, 145]. This value can be obtained by removing one in every four particles from a face-centered cubic (fcc) arrangement, leading to $\eta=\pi/(4\sqrt{2})\approx 0.55536$. The associated space-filling honeycomb (the 3D equivalent of a planar tiling) consists of congruent square bipyramids and is referred to as a square bipyramidal honeycomb or oblate octahedrille, see the ESI [64] for additional information. However, it should be noted that various values of $\eta$ for (frictional) jamming are available, ranging from $\eta\approx 0.54$ [48, 135] to $\eta\approx 0.64$ [146].
It seems likely that the details of the surface interactions and preparation procedure matter. However, it may also be that there are competing regular honeycombs that interfere with obtaining a uniform value across experiments, be those in a lab or using a computer. For example, I obtain a volume fraction of $\eta=8\pi/(27\sqrt{3})\approx 0.537422$, when the spheres are placed using the underlying structure of a (non-gyrated) tetrahedral-octahedral honeycomb. That is, in this arrangement, the particles are the inscribed spheres to the octahedra comprising the honeycomb. This packing fraction is relatively close to the reported RLP value for frictional spheres in 3D, $\eta\approx 0.536$ [48]. Similarly, values of $\eta$ close to where a divergence in the effective viscosity of a sphere suspension with sliding and rolling friction is reported [81], can be obtained by considering other congruent honeycombs. That is, a diamond lattice of triakis truncated tetraheda gives $\eta=\sqrt{3}\pi/16\approx 0.34009$, while Singh et al. report $\eta\approx 0.36$ using their simulations, see the ESI [64].
Turning to the 3D hard-sphere glass transition, the literature places the transition over a wide range of $\eta$ close to $\eta=0.6$. For example, Zaccarelli et al. [147] provide $\eta_{g}\approx 0.588$ using a power-law fit on the MSD the largest particles in a in a polydisperse sample — a value that is also reported for some jamming papers [81, 148] — while an exponential fit leads to a value of $\eta_{g}\approx 0.639$. A recent work by Geiger et al. [149] on a bidisperse sphere mixture places the transition at $\eta_{g}\approx 0.61$. Given the spread, it seems sensible to invert the question of finding a geometric ground state based on a congruent honeycomb for a given volume fraction from the literature. A preliminary search, see the ESI [64], did not reveal any candidate structure that naturally falls into this range. The full exploration of honeycomb structures as potential sources of dynamic arrest is therefore left to follow-up work.